Thomas Jefferson once suggested that the period of a simple pendulum be used to define the...

what would be the period of a pendulum with a length of 1.0m?
Answer #1
T = 2pi sqrt [ 1 / 9.8 ] = 2 s
answered by: candie
Answer #2

Formula is t = 2π x√(l/g) where l = length of 1.0m and g = 9.8

thus 2 x π x √(1/9.8) = 2.006 ≈ 2 seconds

answered by: At A Crossroads
Answer #3
T = 2 pi ( l/g)^0.5

l = 1m eter

Time period = T = 2 seconds
answered by: jazmeen